# NSR-Boost: A Neuro-Symbolic Residual Boosting Framework for Industrial Legacy Models

### v1

@@ Abstract.
Although the Gradient Boosted Decision Trees (GBDTs) dominate industrial tabular applications, upgrading legacy models in high-concurrency production environments still faces prohibitive retraining costs and systemic risks. To address this problem, we present NSR-Boost, a neuro-symbolic residual boosting framework designed specifically for industrial scenarios. Its core advantage lies in being “non-intrusive”. It treats the legacy model as a frozen model and performs targeted repairs on “hard regions” where predictions fail. The framework comprises three key stages: first, finding hard regions through residuals, then generating interpretable experts by generating symbolic code structures using Large Language Model (LLM) and fine-tuning parameters using Bayesian optimization, and finally dynamically integrating experts with legacy model output through a lightweight aggregator. We report on the successful deployment of NSR-Boost within the core financial risk control system at Qfin Holdings. This framework not only significantly outperforms state-of-the-art (SOTA) baselines across six public datasets and one private dataset, more importantly, shows excellent performance gains on real-world online data. In conclusion, it effectively captures long-tail risks missed by traditional models and offers a safe, low-cost evolutionary paradigm for industry.
Neuro-Symbolic AI, Large Language Models, Gradient Boosting, Legacy Model, Interpretability
@@ 1. Introduction
In the fields of computer vision and natural language processing, deep learning has demonstrated exceptional problem-solving capabilities (Noor and Ige, 2025). However, in industrial settings, especially in core business scenarios, tabular data remains the foundational pillar for supporting business operations (Li et al., 2025). In these high-risks, low-latency application environments, Gradient Boosted Decision Trees (GBDTs, such as XGBoost (Chen and Guestrin, 2016) and LightGBM (Ke et al., 2017)) consistently dominate as the de facto industrial standard due to their outstanding interpretability, training efficiency, and inference speed.
@@
Unlike existing methods that treat the LLM as a black-box predictor, NSR-Boost repositions the LLM as a symbolic code generator. We regard the existing GBDT model as a frozen model and utilize LLM to generate expert functions specifically for regions where the legacy model performs poorly. Our framework guides the LLM to uncover complex non-linear interaction logic that is difficult for GDBTs to capture. Crucially, the final output of NSR-Boost consists of executable Python code. This implies that during the actual deployment phase, we completely eliminate the inference cost of LLMs while preserving full logical interpretability.
Specifically, NSR-Boost introduces a bi-level optimization strategy to ensure its stability. In the outer loop, the LLM continuously refines the structure of Symbolic Experts based on “Negative Constraints”. In the inner loop, the LLM autonomously generates the Bayesian search space and executes the optimization scripts to fine-tune the parameters of the generated functions, which effectively addresses the challenge of simultaneous structure and parameter search in symbolic regression. Finally, a context-aware aggregator dynamically integrates the legacy model with the generated experts to ensure a “safe improvement” over the legacy model. Our main contributions are summarized as follows:
@@ 4. Methodology
In this section, we present the overall architecture of NSR-Boost, which is organized into three tightly coupled phases, as illustrated in Figure 1. First, we analyze the residual distribution of the frozen model and employ a decision tree to identify the hard region $\mathcal{D}_{h}$ as the target for subsequent optimization. This ensures that computational resources are concentrated on the model’s weakest points. Subsequently, we initiate $K$ independent chains, adopting a bi-level optimization strategy for expert code generation. Here, we introduce a Cascaded Validation mechanism to ensure that the selected experts contribute substantial metric improvements to the legacy model. Finally, during the inference phase, we deploy a lightweight context-aware aggregator to perform dynamic arbitration on the outputs of each expert, yielding the final prediction.
@@ 4.1. Residual-Guided Region Identification
To precisely map the distribution of these samples within the feature space and construct the partition $\mathbb{P}$ defined in Section 3, we train a shallow CART tree to fit the absolute residuals $|r_{i}|$ (Loh, 2011). We don’t need to be overly precise here, as we will refine the distribution boundaries for it in section 4.2. Subsequently, we introduce a weighted scoring mechanism that calculates a priority score $C$ for each leaf node based on its cumulative error and sample coverage. Leaf nodes with scores exceeding a predefined threshold are identified as “hard regions”. We define the union of samples covered by these selected nodes as the global hard region $\mathcal{D}_{h}$. Crucially, leveraging the inherent disjoint nature of decision tree leaves, $\mathcal{D}_{h}$ is naturally partitioned into multiple disjoint sub-regions, where each region serves as the specific feature subspace for the parallel chain. This strategy effectively balances error magnitude with sample size, guiding the LLM to focus on high-potential regions that offer the greatest leverage for maximizing model performance.
There are several motivations for introducing the CART tree. First, its computational overhead is negligible, imposing no significant training burden. Second, the explicit decision rules derived from the CART tree (e.g., Age > 50) can be embedded into the Prompt as a semantic prior. This distinct boundary information not only assists the LLM in generating highly targeted Symbolic Experts but also effectively mitigates noise interference from expert on samples in non-target regions. Finally, benefiting from the mutual exclusivity of decision tree leaf nodes, the derived regions are mathematically disjoint. This guarantees that the optimization processes across different chains remain non-interfering.
@@ 4.2. Bi-Level Optimization Expert Generation
We decouple the search for discrete symbolic structures from continuous parameter optimization into two nested loops. This strategy stems from the inherent discrete-continuous duality of SR tasks: identifying the optimal functional form is essentially a discrete combinatorial optimization problem, which aligns naturally with the logical reasoning strengths of LLMs (Romera-Paredes et al., 2024). Conversely, determining optimal constant coefficients constitutes a continuous numerical optimization problem demanding high precision, which typically exposes the limitations of token-based prediction in LLMs (Schick et al., 2023). This decoupling effectively mitigates the combinatorial explosion challenge associated with mixed search spaces.
@@
To avoid the syntactic risks introduced by LLM generating complete code, at this stage, the system pre-integrates a standard Tree-structured Parzen Estimator (TPE) optimization framework. The LLM is strictly restricted to the role of a “Configurator”, responsible solely for generating code snippets that define parameters and search spaces (Ma et al., 2024). The prompt are shown in Appendix A.4. These configuration snippets are injected into a sandbox environment for attempted execution. Upon capturing any runtime exceptions, the error stack traces are converted into debugging reports and fed back to the LLM, triggering a targeted correction of the parameter definition logic(Wang et al., 2023). This “Generation-Injection-Error-Correction” closed loop iterates until the script is successfully launched, thereby ensuring the robustness of the automated workflow.
@@
Crucially, to compensate for the discretization error induced by the depth constraints of the decision tree, we explicitly incorporate the hard thresholds of the global conditional statements (generated by the outer loop) into the hyperparameter search space during the first successful Bayesian optimization run. Via numerical iterative approximation, the optimizer is able to fine-tune these coarse-grained boundaries (e.g., adjusting a split point from 50 to 48.5), thereby realizing a seamless transition from discrete symbolic rules to precise numerical boundaries.
@@ 4.3. Context-Aware Aggregation
By training $\mathcal{M}_{g}$, it is capable of dynamically adjusting the contribution weights of each expert based on the input context. Notably, although the input dimension theoretically scales with $K$, thanks to the spatial mutual exclusivity derived from the decision tree partitioning, at most only the local expert covering the specific sample is activated for any given inference query. This sparse activation mechanism ensures that inference latency remains constant, perfectly aligning with strict industrial requirements for low latency.
@@ 5. Experiments
Table 1. Performance comparison on various datasets. The evaluation metric is Accuracy, except for the private dataset (marked with “*”) which is evaluated using AUC. Best results are highlighted in bold, and second-best are underlined. Due to space constraints, detailed standard deviations are provided in Appendix Table 3.
@@ 5.1. Experimental Setup
In our experiments, we utilize Seed-OSS-36B-Instruct as the foundational language model (Team, 2025). We set the sampling temperature to $t=0.1$ and the number of experts to $K=4$. To ensure consistency with baseline methods, NSR-Boost supplies 20 data samples to the LLM during each iteration and terminates after 5 successful iterations, designating the resulting function as the final expert function.
@@ 5.3. Interpretability and Ablation Study
Figure 3. Illustration of the explicit risk rules and actionable high-value segments identified by the Symbolic Expert.
Intrinsic Interpretability and Semantic Reasoning. NSR-Boost provides intrinsic white-box interpretability by generating explicit symbolic expressions. Figure 3 visually illustrates the backbone of a Symbolic Expert trained on the credit-g dataset, alongside its corresponding decision boundary. The internal logic reveals the framework’s unique advantages in semantic reasoning.
First, NSR-Boost demonstrates precise, rule-based subspace localization capabilities. The Symbolic Expert defines a local activation domain via explicit conditional constraints (e.g., checking > 2.16), strictly confining the corrective action to a specific cohort characterized by low checking balances and an age below 35. This mechanism effectively prevents global overfitting during residual learning, ensuring zero interference with non-target groups. Building on this, the automatically constructed variable repay_ease is mathematically equivalent to a logarithmic ratio, while physically corresponding to the inverse of repayment pressure per unit time. This ability to automatically discover high-order features with explicit economic interpretations significantly distinguishes NSR-Boost from the implicit feature extraction typical of traditional neural networks.
Furthermore, our Symbolic Experts capture complex non-monotonic relationships often missed by legacy models through non-linear risk profile modeling. The Gaussian function $\exp(-\frac{(age-\mu)^{2}}{2\sigma^{2}})$ introduced in the code precisely models the non-linear impact of age, identifying a low-risk distribution centered at $27.5$ years and interacting multiplicatively with the repayment pressure feature. As illustrated by the decision surface on the right of Figure 3, this logic successfully corrects the bias inherent in linear models that tend to apply monotonic risk assessments to younger demographics.
@@
Impact of Key Components. Figure 4 presents the ablation study results of NSR-Boost on the credit-g dataset, verifying the necessity of each module.
Firstly, Bayesian optimization holds a central position in the framework. Removing this component has the greatest impact on performance, causing the accuracy to drop from $0.815$ to $0.766$, which clearly demonstrates that although LLM excels at generating symbolic code structures, it is difficult to directly determine precise continuous parameters. It must rely on a hybrid optimization strategy to make up for the shortcomings in numerical accuracy. Secondly, the absence of context-aware aggregation and region identification both lead to an accuracy drop to $0.773$. This confirms that merely generating the correct residual function is not enough. It must be combined with the identification of hard regions and a robust dynamic fusion mechanism to avoid the noise caused by forced fitting by expert models in the entire sample space.
@@ 5.4. Online Deployment
We next report the online deployment results of NSR-Boost on two representative scenarios, where each deployment processes millions of real-world samples.
@@ 5.5. Efficiency and Hyperparameter Analysis
Figure 7. Inference latency vs. accuracy Pareto frontier.
Inference Latency and Pareto Frontier Analysis. Industrial applications impose stringent requirements on response times for online services, typically mandating that single-inference latency be controlled within the millisecond range to support high-concurrency demands. To evaluate the deployment feasibility of NSR-Boost, we compare the trade-off between test accuracy and inference latency across different methods in Figure 7. As illustrated, although deep tabular models such as FT-Transformer and TabNet exhibit competitive accuracy, their inference latencies typically reach the order of 10ms or even 100ms. This prohibitive computational cost not only necessitates expensive GPU clusters but also fails to meet the rigid low-latency benchmarks required for real-time risk control scenarios. In contrast, NSR-Boost is positioned on the Pareto Frontier of the accuracy-latency trade-off. By leveraging its generated pure Python code format, NSR-Boost maintains SOTA accuracy while constraining inference latency to under 1ms. Compared to OpenFE, we achieve higher accuracy at equivalent latency levels; compared to XGBoost, we realize a significant leap in accuracy with only a negligible increase in computational overhead (at the microsecond level). This demonstrates that NSR-Boost is currently the unique solution capable of simultaneously satisfying the dual criteria of high precision and low latency, enabling seamless deployment without requiring upgrades to existing hardware infrastructure.
Figure 8. Impact of the number of Symbolic Expert on the industrial financial private dataset.
Hyperparameter Sensitivity and Residual Mechanism Verification. To identify the optimal balance between inference precision and system load for industrial deployment, we conducted a sensitivity analysis on the number of the Symbolic Expert ($K$) using a large-scale private dataset. As illustrated in the left panel of Figure 8, the AUC of the hybrid model exhibits a typical trend of diminishing returns as the number of experts increases. The curve displays a distinct “Knee Point” at $K=4$: at this juncture, the model has already secured the vast majority of performance gains. Continuing to increase the number of experts yields only negligible improvements in accuracy while linearly increasing inference latency.
@@ A.2. Dataset Details
Table 2 details the collection of binary classification datasets utilized for evaluation, primarily sourced from OpenML, alongside a private dataset from Qfin Holdings. Given that the high-stakes domains discussed in this paper are typically modeled as binary classification problems, we place particular emphasis on such tasks. To align with business requirements for interpretability, we specifically selected datasets possessing descriptive feature names. Each dataset is accompanied by a task description to facilitate an understanding of the background context. Our selection spans from small-scale to large-scale datasets. This diverse setting, characterized by varying feature dimensionalities and sample sizes, represents a wide spectrum of real-world scenarios, thereby providing a robust framework for evaluating the proposed method.
Table 2. Dataset statistics.
@@
private-data 200 497677 Qfin -

### v2

@@ Abstract.
Although the Gradient Boosted Decision Trees (GBDTs) dominate industrial tabular applications, upgrading legacy models in high-concurrency production environments still faces prohibitive retraining costs and systemic risks. To address this problem, we present NSR-Boost, a neuro-symbolic residual boosting framework designed specifically for industrial scenarios. Its core advantage lies in being “non-intrusive”. It treats the legacy model as a frozen model and performs targeted repairs on “hard regions” where predictions fail. The framework comprises three key stages: First, finding hard regions through residuals, then generating interpretable experts by generating symbolic code structures using Large Language Model (LLM) and fine-tuning parameters using Bayesian optimization, and finally dynamically integrating experts with legacy model output through a lightweight aggregator. Experimental results demonstrate that the framework not only significantly outperforms state-of-the-art (SOTA) baselines across six public datasets and one private dataset. More importantly, we report the successful deployment of NSR-Boost within the core financial risk control system of Qfin Holdings, where empirical results on real-world online traffic exhibit superior performance improvements and a significant reduction in the bad rate. In conclusion, it effectively captures long-tail risks missed by traditional models and offers a safe, low-cost evolutionary paradigm for industry.
Neuro-Symbolic AI, Large Language Models, Tabular Data, Legacy Model, Interpretability
@@ 1. Introduction
In the fields of computer vision and natural language processing, deep learning has demonstrated exceptional problem-solving capabilities (Noor and Ige, 2025). However, in industrial settings, especially in core business scenarios, tabular data remains the foundational pillar for supporting business operations (Li et al., 2025). In these high-risk, low-latency application environments, Gradient Boosted Decision Trees (GBDTs, such as XGBoost (Chen and Guestrin, 2016) and LightGBM (Ke et al., 2017)) consistently dominate as the de facto industrial standard due to their outstanding interpretability, training efficiency, and inference speed.
@@
Unlike existing methods that treat the LLM as a black-box predictor, NSR-Boost repositions the LLM as a symbolic code generator. We regard the existing GBDT model as a frozen model and utilize LLM to generate expert functions specifically for regions where the legacy model performs poorly. Our framework guides the LLM to uncover complex non-linear interaction logic that is difficult for GBDTs to capture. Crucially, the final output of NSR-Boost consists of executable Python code. This implies that during the actual deployment phase, we completely eliminate the inference cost of LLMs while preserving full logical interpretability.
Specifically, NSR-Boost introduces a bi-level optimization strategy to ensure its stability. In the outer loop, the LLM continuously refines the structure of Symbolic Experts based on “Negative/Positive Constraints”. In the inner loop, the LLM autonomously generates the Bayesian search space and executes the optimization scripts to fine-tune the parameters of the generated functions, which effectively addresses the challenge of simultaneous structure and parameter search in symbolic regression. Finally, a context-aware aggregator dynamically integrates the legacy model with the generated experts to ensure a “safe improvement” over the legacy model. Our main contributions are summarized as follows:
@@ 4. Methodology
In this section, we present the overall architecture of NSR-Boost, which is organized into three tightly coupled phases, as illustrated in Figure 1. First, we analyze the residual distribution of the frozen model and employ a CART tree to identify the hard region $\mathcal{D}_{h}$ as the target for subsequent optimization. This ensures that computational resources are concentrated on the model’s weakest points. Subsequently, we initiate $K$ independent chains, adopting a bi-level optimization strategy for expert code generation. Here, we introduce a Cascaded Validation mechanism to ensure that the selected experts contribute substantial metric improvements to the legacy model. Finally, during the inference phase, we deploy a lightweight context-aware aggregator to perform dynamic arbitration on the outputs of each expert, yielding the final prediction.
@@ 4.1. Residual-Guided Region Identification
To precisely map the distribution of these samples within the feature space and construct the partition $\mathbb{P}$ defined in Section 3, we train a shallow CART tree to fit the absolute residuals $|r_{i}|$ (Loh, 2011). We don’t need to be overly precise here, as we will refine the distribution boundaries for it in section 4.2. Subsequently, we introduce a weighted scoring mechanism that calculates a priority score $C$ for each leaf node based on its cumulative error and sample coverage. Leaf nodes with scores exceeding a predefined threshold are identified as “hard regions”. We define the union of samples covered by these selected nodes as the global hard region $\mathcal{D}_{h}$. Crucially, leveraging the inherent disjoint nature of CART tree leaves, $\mathcal{D}_{h}$ is naturally partitioned into multiple disjoint sub-regions, where each region serves as the specific feature subspace for the parallel chain. This strategy effectively balances error magnitude with sample size, guiding the LLM to focus on high-potential regions that offer the greatest leverage for maximizing model performance.
There are several motivations for introducing the CART tree. First, its computational overhead is negligible, imposing no significant training burden. Second, the explicit decision rules derived from the CART tree (e.g., Age > 50) can be embedded into the Prompt as a semantic prior. This distinct boundary information not only assists the LLM in generating highly targeted Symbolic Experts but also effectively mitigates noise interference from expert on samples in non-target regions. Finally, benefiting from the mutual exclusivity of CART tree leaf nodes, the derived regions are mathematically disjoint. This guarantees that the optimization processes across different chains remain non-interfering.
@@ 4.2. Bi-Level Optimization Expert Generation
We decouple the search for discrete symbolic structures from continuous parameter optimization into two nested loops. This strategy stems from the inherent discrete-continuous duality of SR tasks: identifying the optimal functional form is essentially a discrete combinatorial optimization problem, which aligns naturally with the logical reasoning strengths of LLMs (Romera-Paredes et al., 2024); conversely, determining optimal constant coefficients constitutes a continuous numerical optimization problem demanding high precision, which typically exposes the limitations of token-based prediction in LLMs (Schick et al., 2023). This decoupling effectively mitigates the combinatorial explosion challenge associated with mixed search spaces.
@@
To avoid the syntactic risks introduced by LLM generating complete code, at this stage, the system pre-integrates a standard Tree-structured Parzen Estimator (TPE) optimization framework. The LLM is strictly restricted to the role of a “Configurator”, responsible solely for generating code snippets that define parameters and search spaces (Ma et al., 2024). The prompt is shown in Appendix A.4. These configuration snippets are injected into a sandbox environment for attempted execution. Upon capturing any runtime exceptions, the error stack traces are converted into debugging reports and fed back to the LLM, triggering a targeted correction of the parameter definition logic(Wang et al., 2023). This “Generation-Injection-Error-Correction” closed loop iterates until the script is successfully launched, thereby ensuring the robustness of the automated workflow.
@@
Crucially, to compensate for the discretization error induced by the depth constraints of the CART tree, a hard threshold of global conditional statements generated by the outer loop is explicitly incorporated into the hyperparameter search space during the first successful Bayesian optimization run. Via numerical iterative approximation, the optimizer is able to fine-tune these coarse-grained boundaries, for example, adjusting a split point from 50 to 48.5, thereby realizing a seamless transition from discrete symbolic rules to precise numerical boundaries.
@@ 4.3. Context-Aware Aggregation
By training $\mathcal{M}_{g}$, it is capable of dynamically adjusting the contribution weights of each expert based on the input context. Notably, although the input dimension theoretically scales with $K$, thanks to the spatial mutual exclusivity derived from the CART tree partitioning, at most only the local expert covering the specific sample is activated for any given inference query. This sparse activation mechanism ensures that inference latency remains constant, perfectly aligning with strict industrial requirements for low latency.
@@ 5. Experiments
Table 1. Performance comparison on various datasets. The evaluation metric is accuracy, except for the private dataset (marked with $\dagger$”) which is evaluated using AUC. Best results are highlighted in bold, and second-best are underlined. Due to space constraints, detailed standard deviations are provided in Appendix Table 3.
@@ 5.1. Experimental Setup
In our experiments, we employ the locally deployed Seed-OSS-36B-Instruct (Team, 2025a) to ensure strict data compliance while leveraging raw feature semantics. We set the sampling temperature to $t=0.1$ and the number of experts to $K=5$. To ensure consistency with baseline methods, NSR-Boost supplies 20 data samples to the LLM during each iteration and terminates after 5 successful iterations, designating the resulting function as the final expert function.
@@ 5.3. Interpretability and Ablation Study
Figure 3. The interpretable logic of a Symbolic Expert.
Figure 3 illustrates the generated symbolic expert code generated on the credit-g dataset and intuitively demonstrates the white-box interpretability and safety mechanisms of NSR-Boost. The code employs explicit if-else logic structures to precisely delineate the operational boundaries of residual correction. The model activates and computes additional correction terms only when samples fall within specific regions. It directly outputs the original predictions of the legacy model in all other instances to ensure zero interference with non-target groups. Furthermore the automatically constructed features possess clear business semantics that enable business personnel to directly review and comprehend the decision logic. This implementation thereby achieves a unification of high performance and high transparency. A more detailed analysis is provided in Appendix B.2.
@@
Figure 4 presents the ablation study results of NSR-Boost on the credit-g dataset, verifying the necessity of each module. Firstly, Bayesian optimization holds a central position in the framework. Removing this component has the greatest impact on performance, causing the accuracy to drop from $0.815$ to $0.766$, which clearly demonstrates that although LLM excels at generating symbolic code structures, it is difficult to directly determine precise continuous parameters. It must rely on a hybrid optimization strategy to make up for the shortcomings in numerical accuracy. Secondly, the absence of context-aware aggregation and region identification both lead to an accuracy drop to $0.773$. This confirms that merely generating the correct residual function is not enough. It must be combined with the identification of hard regions and a robust dynamic fusion mechanism to avoid the noise caused by forced fitting by expert models in the entire sample space.
@@ 5.4. Online Deployment
Next, we report the online deployment results of NSR-Boost in two typical scenarios, each processing millions of real-world samples. Since NSR-Boost operates as an additional enhancement to the legacy model, the prediction results of both the legacy model and NSR-Boost are kept and recorded during the online deployment process, enabling a direct comparison of the two on the same input request without traffic splitting.
@@ 5.5. Efficiency and Hyperparameter Analysis
Figure 7. Inference latency vs. accuracy Pareto frontier for different methods.
Inference Latency and Pareto Frontier Analysis. Industrial applications impose stringent requirements on response times for online services, typically mandating that single-inference latency be controlled within the millisecond range to support high-concurrency demands. To evaluate the deployment feasibility of NSR-Boost, we compare the trade-off between test accuracy and inference latency across different methods in Figure 7. As illustrated, although deep tabular models such as FT-Transformer and TabNet exhibit competitive accuracy, their inference latencies typically reach the order of 10ms or even 100ms. This prohibitive computational cost not only necessitates expensive GPU clusters but also fails to meet the rigid low-latency benchmarks required for real-time risk control scenarios. In contrast, NSR-Boost is positioned on the Pareto Frontier of the accuracy-latency trade-off. By leveraging its generated pure Python code format, NSR-Boost maintains SOTA accuracy while constraining inference latency to under 1ms. Compared to OpenFE, we achieve higher accuracy at equivalent latency levels. Compared to XGBoost, we realize a significant leap in accuracy with only a negligible increase in computational overhead (at the microsecond level). This demonstrates that NSR-Boost is currently the unique solution capable of simultaneously satisfying the dual criteria of high precision and low latency, enabling seamless deployment without requiring upgrades to existing hardware infrastructure.
Figure 8. Impact of the number of Symbolic Experts on the industrial financial private dataset.
Hyperparameter Sensitivity and Residual Mechanism Verification. To identify the optimal balance between inference precision and system load for industrial deployment, we conducted a sensitivity analysis on the number of Symbolic Experts ($K$) using a large-scale private dataset. As illustrated in the left panel of Figure 8, the AUC of the hybrid model exhibits a typical trend of diminishing returns as the number of experts increases. The curve displays a distinct “Knee Point” at $K=4$: At this juncture, the model has already secured the vast majority of performance gains. Continuing to increase the number of experts yields only negligible improvements in accuracy while linearly increasing inference latency.
@@ A.2. Dataset Details
Table 2 details the collection of binary classification datasets utilized for evaluation, primarily sourced from OpenML, alongside a private dataset from Qfin Holdings. Given that the high-stakes domains discussed in this paper are typically modeled as binary classification problems, we place particular emphasis on such tasks. In order to meet the interpretability requirements of the business, we deliberately choose those datasets with descriptive feature names. Each dataset is accompanied by a task description to facilitate an understanding of the background context. Our selection spans from small-scale to large-scale datasets. This diverse setting, characterized by varying feature dimensionalities and sample sizes, represents a wide spectrum of real-world scenarios, thereby providing a robust framework for evaluating the proposed method.
@@
private-data 200 497677 Qfin Holdings -
Table 2. Dataset statistics.